1 Exercises

Exercise 1 An exercise with a solution. [Solution] Exercise 2 An exercise with a hint and a secret solution. [Hint] [Homework] Exercise 3 An exercise with a hint. [Hint]

2 test: Solutions

2.1 Some Solutions

Solution 1 This is a solution. Hint 2 This is a hint. Hint 3 This is a hint.

2.2 Extra Solutions

Solution to 2 This is a secret solution.